# Viscous-inertial waves on the surface of the Sun: modeling, forward and inverse problems

### v2

@@ 3.4 Additional regularity
$f_{m}\in L^{2}({r{\mathbb{S}^{2}}};m),\qquad\Omega\in H^{2}({r{\mathbb{S}^{2}}};m=0).$ (26)
@@
$\begin{aligned} \forall\psi_{m}\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m):\quad&\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle=\left\langle\tilde{f}_{m},\psi_{m}\right\rangle\\ &\mbox{with }\tilde{f}_{m}:=f_{m}+i\omega\Delta_{m}\Psi_{m}-im\beta_{\Omega}\Delta_{m}\Psi_{m}+im\alpha_{\Omega}\Psi_{m}.\end{aligned}$ (28)
@@
It follows from the isometry property (62) that weak solutions $\Psi\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ to the biharmonic equation $\Delta_{\mathrm{h}}^{2}\Psi=f$ with $f\in L^{2}({r{\mathbb{S}^{2}}})$ belong to $H^{4}({r{\mathbb{S}^{2}}})$. Due to the separability of $\Delta_{\mathrm{h}}$, weak solutions $\Psi_{m}\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ to the separated equation $\Delta_{m}^{2}\Psi_{m}=\tilde{f}_{m}$ with $\tilde{f}_{m}\in L^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ also belong to $H^{4}_{\diamond}({r{\mathbb{S}^{2}}};m)$ with $\|\Psi_{m}\|_{H^{4}}=\|\tilde{f}_{m}\|_{L^{2}}$. Together with $\|\Psi_{m}\|_{H^{2}}\leq\|\mathbf{B}_{m}^{-1}\|\|f_{m}\|_{H^{-2}}\leq\|\mathbf{B}_{m}^{-1}\|\|f_{m}\|_{L^{2}}$ this shows that $\|\Psi_{m}\|_{H^{4}}\leq C\|f\|_{L^{2}}$, i.e., $\mathbf{B}_{m}:H^{4}_{\diamond}({r{\mathbb{S}^{2}}};m)\to L^{2}({r{\mathbb{S}^{2}}};m)$ is boundedly invertible. The proof of the analogous statement for $\mathbf{B}_{m}^{\star}$ only requires a different treatment of the term involving $\beta_{\Omega}$. Here, we use the fact that $H^{2}({r{\mathbb{S}^{2}}})$ is a Banach algebra (see [2, p. 115] or [3, Thm 1.4]) to show that $\|\beta_{\Omega}\Psi_{m}\|_{H^{2}}\leq C\|\beta_{\Omega}\|_{H^{2}}\|\Psi_{m}\|_{H^{2}}$.
@@ 3.5 Well-posedness on the restricted domain
Inspired by the practical measurement scenario in Section 2.3, where only data on the restricted latitude range $I^{\prime}=(\epsilon,\pi-\epsilon)$ can be observed, this section derives well-posedness, regularity of the wave operators on this restricted domain. This result will be later employed in Section 5.2 for convergence guarantee of reconstruction method under restricted measurements.
We sketch how an analysis similar to that of Propositions 7 and 8 can be carried out, whose key elements are symmetry of the bi-Laplacian and Laplacian operators and partial integrations. According to Remark 20, symmetry of Laplacian operators on function spaces supported on $I^{\prime}$ is guaranteed with homogeneous Cauchy boundary data, i.e. $\Gamma_{2}$ in (10).
To this end, we define the following function spaces on $I^{\prime}$ with norms inherited from $H^{s}({r{\mathbb{S}^{2}}};m)$:
@@
$\Psi|_{\partial I^{\prime}}=u_{1}\in\mathbb{C}^{2}\quad\text{and}\quad\Psi^{\prime}|_{\partial I^{\prime}}=u_{2}\in\mathbb{C}^{2}.$ (31)
@@
Concerning the lifted regularity result as in Proposition 8, by setting the auxiliary source $\tilde{f}_{m}$ as in (28) and partial integration, we (formally) have $\gamma\left\langle\Delta_{m}^{2}\Psi_{m},\psi_{m}\right\rangle=\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle=\left\langle\tilde{f}_{m},\psi_{m}\right\rangle$ for all $\psi_{m}\in H^{2}_{0}(I^{\prime};m)$. Using the fact the weak solutions of the bi-Laplace equations are also strong solutions, one obtains boundedness of the fourth order term $\|\Delta_{m}^{2}\Psi_{m}\|_{L^{2}}$, given higher regularity of $\Omega$ and $f$ as in Proposition 8. Alternatively, one can apply general elliptic regularity results as, e.g., in [38, Sec. 5.3.4]. ∎
4 Inverse problems for viscosity and rotation
@@ 4.1 Inversion with different observation strategies
In this section, we address the inverse problem of recovering the differential rotation function $\Omega$ and the viscosity coefficient $\gamma$ in the separated equations. To achieve this, we require additional information, namely full or partial measurements of the state, as described in Section 2.3:
$\begin{array}[]{l}L:V_{m}\to Y\\ (L\Psi)(\theta):=\Psi(\theta)\end{array}\qquad Y=\begin{cases}L^{2}(I,r^{2}\sin),&\text{full measurements}\\ L^{2}(I^{\prime},r^{2}\sin),&\text{partial measurements}\end{cases}$ (34)
with the weighted $L^{2}$-data spaces. We formulate the parameter identification problem within the Hilbert space framework established by the well-posedness results in Sections 3.2-3.3:
$\gamma\in\mathbb{R}^{+},\quad\Omega\in X:=H^{1}({r{\mathbb{S}^{2}}},m=0),\quad\Psi\in V_{m}:=H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m),\quad y\in Y,$ (35)
By introducing the parameter-to-state map
@@
together with the measurement operator we can formulate the forward operator $F$ for the parameter identification as
$F:\mathcal{D}(S)\to Y\qquad F:=L\circ S.$ (37)
Note that $X$ is the space of real-valued functions, while $V_{m}$ and $Y$ are spaces of complex-valued functions; together, they form Gelfand triples.
@@ 4.2 Sensitivity and adjoints
Compactness of the embedding $V_{m}\hookrightarrow L^{2}(I,r^{2}\sin)$ yields compactness of the measurement operator $L$, thus of the forward operator $F$. In such cases, the inverse coefficient problem becomes ill-posed, necessitating regularization for stable reconstruction. Given the inherent nonlinearity of the problem, iterative methods such as Landweber-type and Newton-type algorithms are typically employed for reconstruction (see Section 5). For noisy data $y^{\delta}\in Y$, the former is given by
$p_{k+1}=p_{k}-F^{\prime}(p_{k})^{*}(F(p_{k})-y^{\delta})\qquad k\leq K(y^{\delta},\delta),$
where the unknown parameter is $p:=(\gamma,\Omega)$. The stopping index $K(y^{\delta},\delta)$ serves as the regularization parameter and is determined according to the discrepancy principle; see Section 6 for further details. This formulation highlights that the main components of iterative gradient-based methods: sensitivity analysis and the derivation of the adjoint; these are the primary focus of this section.
To this end, we will add the dependence on the unknown parameters $p=(\gamma,\Omega)\in\mathcal{D}(S)$ in parenthesis, i.e. $\mathbf{B}_{m}(p)=\mathbf{B}_{m}:V_{m}\to V_{m}^{*}.$ Since the mapping $p\mapsto\mathbf{B}_{m}(p)$ is affine linear, its derivative with respect to $p$ is independent of $p$ and we can omit it in our notation. Hence, we simply denote the derivative of $p\mapsto\mathbf{B}_{m}(p)$ by
$\mathbf{B}_{m}^{\prime}:\mathbb{R}\times X\to\mathcal{L}(V_{m},V_{m}^{*}),\qquad\mathbf{B}_{m}^{\prime}(\delta\gamma,\delta\Omega)\Psi:=(\delta\gamma)\Delta_{m}^{2}\Psi+im\alpha_{\delta\Omega}\Psi-im(\delta\Omega)\Psi.$
@@
The result follows from the differentiability statement in the implicit function theorem (see, e.g., [41, §4.7]) applied to the operator $G:\mathcal{D}(S)\times V_{m}\to V_{m}^{*}$, $G(p,\Psi):=\mathbf{B}_{m}(p)\Psi-f$. Here, one interprets $S$ as the implicitly defined function $G(p,S(p))=0$ and employs linearity of $G$ in the second argument as well as bounded invertibility of $\mathbf{B}_{m}(p):V_{m}\to V_{m}^{*}$. ∎
Differentiability of $F$ is straightforward from that of $S$ and boundedness of the linear measurement operator $L$ defined in (34). We now derive the adjoint for different observation senarios.
@@
$F^{\prime}[p]^{*}:Y\to\mathbb{R}\times X\qquad F^{\prime}[p]^{*}y=-\left[\begin{smallmatrix}1&0\\ 0&I_{X}\end{smallmatrix}\right]\,\Re\left([\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}(\mathbf{B}_{m}(p)^{-1})^{\star}y\right)$ (39)
@@
$[\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}:V_{m}\to\mathbb{R}\times X^{*},\quad(\mathbf{B}_{m}(p)^{-1})^{\star}:V_{m}^{*}\to V_{m}$ (40)
@@
$\quad=\int_{0}^{\pi}\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}\left(\delta\Omega\sin^{2}\theta\right)\right)\Psi\overline{z}\,d\theta=\int_{0}^{\pi}\frac{-1}{\sin\theta}\frac{d}{d\theta}\left(\delta\Omega\sin^{2}\theta\right)\frac{d}{d\theta}\left(\Psi\overline{z}\right)\,d\theta$
$\quad=\int_{0}^{\pi}(\delta\Omega)(\theta)\sin^{2}\theta\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}\left(\Psi\overline{z}\right)\right)\,d\theta=\left\langle\delta\Omega\,,\,\frac{\sin}{r^{2}}\frac{d}{d\theta}\left(\frac{1}{\sin}\frac{d}{d\theta}\left(\overline{\Psi}z\right)\right)\right\rangle$
@@
$R^{*}:L^{2}(I^{\prime},r^{2}\sin)\to L^{2}(I,r^{2}\sin)\qquad R^{*}\Psi)(\theta):=\begin{cases}\Psi(\theta),&\theta\in I^{\prime},\\ 0,&\theta\in I\setminus I^{\prime}.\end{cases}$
The corresponding forward operator relates to the forward operator with full measurement via $F_{\mathrm{part}}[p]=RF_{\mathrm{full}}[p]$. Thus, $F_{\mathrm{part}}^{\prime}[p]^{*}=F_{\mathrm{full}}^{\prime}[p]^{*}R^{*}$ with $F_{\mathrm{full}}^{\prime}[p]^{*}$ as in Corollary 12.
@@ 5 Tangential cone condition
As discussed in Section 4, the inverse problem under consideration is ill-posed, necessitating the use of regularization to achieve stable reconstruction of the target coefficients. Iterative gradient-based regularization methods, such as Landweber iteration, Newton-type schemes, and their variants, fundamentally rely on three core elements: the well-posedness of the forward operator $F$ (Section 3), adjoint of the linearized operator (Section 4.2), and guarantees of convergence. This final requirement is typically ensured by making specific structural assumptions about $F$, particularly regarding its nonlinearity and the uniform boundedness of its derivative; these issues are the main focus of this section.
To address this, we consider the tangential cone condition (TCC), a celebrated criterion first introduced in [16] that ensures local convergence of iterative regularization algorithms. Intuitively, if the forward map $F$ is excessively nonlinear, there is no general guarantee that gradient descent steps will remain within the vicinity of the true solution. The tangential cone condition offers a quantitative tool for assessing the nonlinearity of $F$, especially in the context of compact operators, thereby paving the way for the successful application of iterative regularization methods. Notably, satisfaction of this condition also implies a characterization of the solution set via null space of the linearized operator, uniqueness of the minimum-norm solution, and, combined with the null space condition, ensures that the iterative reconstruction sequence converges to this solution [22].
For a general model $F:X\ni p\mapsto y\in Y$, TCC is stated in the following (strong) form, that is the inequality
@@
holds for some constant $C_{tc}>0$ and radius $R>0$ of the ball around a ground truth $p^{\dagger}\in\mathcal{D}(F)$; see also [25] for the weak forms. In our setting, TCC reads as
@@
It is worth noting that, in addition to the TCC, several other conditions can be employed to characterize the structure of the forward map and provide convergence guarantees. Notable examples include range-invariance [24], convexity, and the Polyak-Lojasiewicz condition [31].
Although the tangential cone condition is a powerful tool, it can be challenging to verify in practice; for a general verification strategy with full data, we refer the reader to [23]. Contributions in this area span a wide range of applications, including elliptic inverse problems [18], elastography [21, 20], electrical impedance tomography [26], full waveform inversion [10], and more recently, neural network-based inverse problems [35, 1].
In this work, we contribute to the literature by establishing the tangential cone condition for full measurements, introducing a lifted regularity strategy developed in Section 3.4 to account for realistic measured data in $L^{2}$. Furthermore, we extend our verification to the setting of restricted measurements.

### v3

@@ 3.4 Additional regularity
We have established unique existence of wave solutions. However, we proceed further with proving higher regularity of the wave solutions, particularly for the separated equations. This regularity result is not merely a technical refinement; it is essential for convergence guarantee of iterative regularization methods later in Section 5.
@@
$f_{m}\in L^{2}({r{\mathbb{S}^{2}}};m),\hskip 18.49988pt\Omega\in H^{2}({r{\mathbb{S}^{2}}};m=0).$ (29)
@@
$\begin{aligned} \forall\psi_{m}\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m):\qquad&\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle=\left\langle\tilde{f}_{m},\psi_{m}\right\rangle\\ &\mbox{with }\tilde{f}_{m}:=f_{m}+i\omega\Delta_{m}\Psi_{m}-im\beta_{\Omega}\Delta_{m}\Psi_{m}+im\alpha_{\Omega}\Psi_{m}.\end{aligned}$ (31)
@@
It follows from the isometry property (65) that weak solutions $\Psi\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ to the biharmonic equation $\Delta_{\mathrm{h}}^{2}\Psi=f$ with $f\in L^{2}({r{\mathbb{S}^{2}}})$ belong to $H^{4}({r{\mathbb{S}^{2}}})$. Due to the separability of $\Delta_{\mathrm{h}}$, weak solutions $\Psi_{m}\in H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ to the separated equation $\Delta_{m}^{2}\Psi_{m}=\tilde{f}_{m}$ with $\tilde{f}_{m}\in L^{2}_{\diamond}({r{\mathbb{S}^{2}}})$ also belong to $H^{4}_{\diamond}({r{\mathbb{S}^{2}}};m)$ with $\|\Psi_{m}\|_{H^{4}}=\|\tilde{f}_{m}\|_{L^{2}}$. Together with $\|\Psi_{m}\|_{H^{2}}\leq\|\mathbf{B}_{m}^{-1}\|\|f_{m}\|_{H^{-2}}\leq\|\mathbf{B}_{m}^{-1}\|\|f_{m}\|_{L^{2}}$ this shows that $\|\Psi_{m}\|_{H^{4}}\leq C\|f\|_{L^{2}}$, i.e., $\mathbf{B}_{m}:H^{4}_{\diamond}({r{\mathbb{S}^{2}}};m)\to L^{2}({r{\mathbb{S}^{2}}};m)$ is boundedly invertible. The proof of the analogous statement for $\mathbf{B}_{m}^{\star}$ only requires a different treatment of the term involving $\beta_{\Omega}$. Here, we use the fact that $H^{2}({r{\mathbb{S}^{2}}})$ is a Banach algebra (see [adams:75, p. 115] or [BadrBernicotRuss, Thm 1.4]) to show that $\|\beta_{\Omega}\Psi_{m}\|_{H^{2}}\leq C\|\beta_{\Omega}\|_{H^{2}}\|\Psi_{m}\|_{H^{2}}$.
@@ 3.5 Well-posedness on the restricted domain
Motivated by the practical measurement scenario in Section 2.3, where observations are limited to a restricted latitude range $I^{\prime}=(\epsilon,\pi-\epsilon)$, this section derives well-posedness and regularity of the wave operators on this restricted domain. These results are subsequently employed in Section 5.2 for convergence analysis of reconstruction methods under restricted observations.
We outline how an analysis analogous to that of Propositions 7 and 8 can be conducted. The key elements are symmetry of the bi-Laplacian and Laplacian operators and strategic use of partial integrations. According to Remark 20, symmetry of Laplacian operators on function spaces supported on $I^{\prime}$ is guaranteed with homogeneous Cauchy boundary data, i.e. $\Gamma_{2}$ in (13). To this end, we define the following function spaces on $I^{\prime}$ with norms inherited from $H^{s}({r{\mathbb{S}^{2}}};m)$:
@@
$\Psi|_{\partial I^{\prime}}=u_{1}\in\mathbb{C}^{2}\qquad\text{and}\qquad\Psi^{\prime}|_{\partial I^{\prime}}=u_{2}\in\mathbb{C}^{2}.$ (34)
@@
Concerning the lifted regularity result as in Proposition 8, by setting the auxiliary source $\tilde{f}_{m}$ as in (31) and partial integration, we (formally) have $\gamma\left\langle\Delta_{m}^{2}\Psi_{m},\psi_{m}\right\rangle=\gamma\left\langle\Delta_{m}\Psi_{m},\Delta_{m}\psi_{m}\right\rangle=\left\langle\tilde{f}_{m},\psi_{m}\right\rangle$ for all $\psi_{m}\in H^{2}_{0}(I^{\prime};m)$. Using the fact the weak solutions of the bi-Laplace equations are also strong solutions, one obtains boundedness of the fourth order term $\|\Delta_{m}^{2}\Psi_{m}\|_{L^{2}}$, given higher regularity of $\Omega$ and $f$ as in Proposition 8. Alternatively, one can apply general elliptic regularity results as, e.g., in [Triebel, Sec. 5.3.4]. ∎
4 Inverse problem for viscosity and rotation
With the forward problems established, we now address the inverse problem: recovering the differential rotation function $\Omega$ and the viscosity parameter $\gamma$ in the separated equations under various data measurement strategies.
@@ 4.1 Inversion with different observation strategies
As described in Section 2.3, data is acquired under two scenarios: full or partial measurements of the wave solution. The measurement operator is hence defined as
$\begin{array}[]{l}L:V_{m}\to Y\\ (L\Psi)(\theta):=\Psi(\theta)\end{array}\hskip 18.49988ptY=\begin{cases}L^{2}(I,r^{2}\sin),&\text{full measurements}\\ L^{2}(I^{\prime},r^{2}\sin),&\text{partial measurements}\end{cases}$ (37)
where $Y$ are weighted $L^{2}$-data spaces. Within the Hilbert space framework established by the well-posedness results in Sections 3.23.3, that is
$\gamma\in\mathbb{R}^{+},\qquad\Omega\in X:=H^{1}({r{\mathbb{S}^{2}}},m=0),\qquad\Psi\in V_{m}:=H^{2}_{\diamond}({r{\mathbb{S}^{2}}};m),\qquad y\in Y,$ (38)
we introduce the parameter-to-state map
@@
Together with the measurement operator $L$ in (37), the forward operator $F$ for the parameter identification is formulated as
$F:\mathcal{D}(S)\to Y\hskip 18.49988ptF:=L\circ S.$ (40)
Note that $X$ is the space of real-valued functions, while $V_{m}$ and $Y$ are spaces of complex-valued functions; together, they form Gelfand triples. Compactness of the embedding $V_{m}\hookrightarrow Y$ yields compactness of the measurement operator $L$, thus of the forward operator $F$. In such cases, the inverse parameter problem becomes ill-posed, necessitating regularization to ensure stable reconstruction of the target coefficients.
@@ 4.2 Sensitivity and adjoints
Given the inherent nonlinearity of the inverse problem, iterative regularization methods such as Landweber-type and Newton-type algorithms are typically employed (see Section 5). For noisy data $y^{\delta}\in Y$, the Landweber iteration is given as
$p_{k+1}=p_{k}-F^{\prime}(p_{k})^{*}(F(p_{k})-y^{\delta})\hskip 18.49988ptk\leq K(y^{\delta},\delta),$
for the unknown parameter $p:=(\gamma,\Omega)$. The stopping index $K(y^{\delta},\delta)$ serves as the regularization parameter and is determined according to the discrepancy principle; see Section 6 for further details. This formulation highlights the main components of gradient-based regularization schemes: sensitivity analysis and adjoint derivation; these are the main focus of this section.
For clarity, we will explicitly denote the dependence of $\mathbf{B}_{m}$ on the unknown parameters $p=(\gamma,\Omega)\in\mathcal{D}(S)$, writing $\mathbf{B}_{m}(p):V_{m}\to V_{m}^{*}.$ As the mapping $p\mapsto\mathbf{B}_{m}(p)$ is affine linear, the derivative is independent of $p$; we thus suppress this argument in the derivative and write
$\mathbf{B}_{m}^{\prime}:\mathbb{R}\times X\to\mathcal{L}(V_{m},V_{m}^{*}),\hskip 18.49988pt\mathbf{B}_{m}^{\prime}(\delta\gamma,\delta\Omega)\Psi:=(\delta\gamma)\Delta_{m}^{2}\Psi+im\alpha_{\delta\Omega}\Psi-im(\delta\Omega)\Psi.$
@@
The result follows from the differentiability statement in the implicit function theorem (see, e.g., [zeidler1986, §4.7]) applied to the operator $G:\mathcal{D}(S)\times V_{m}\to V_{m}^{*}$, $G(p,\Psi):=\mathbf{B}_{m}(p)\Psi-f$. Here, one interprets $S$ as the implicitly defined function $G(p,S(p))=0$ and employs linearity of $G$ in the second argument as well as bounded invertibility of $\mathbf{B}_{m}(p):V_{m}\to V_{m}^{*}$. ∎
Differentiability of $F$ is straightforward from that of $S$ and boundedness of the linear measurement operator $L$. We now derive the adjoint for different observations.
@@
$F^{\prime}[p]^{*}:Y\to\mathbb{R}\times X\hskip 18.49988ptF^{\prime}[p]^{*}y=-\left[\begin{smallmatrix}1&0\\ 0&I_{X}\end{smallmatrix}\right]\,\Re\left([\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}(\mathbf{B}_{m}(p)^{-1})^{\star}y\right)$ (42)
@@
$[\mathbf{B}_{m}^{\prime}(\cdot)\Psi]^{\star}:V_{m}\to\mathbb{R}\times X^{*},\qquad(\mathbf{B}_{m}(p)^{-1})^{\star}:V_{m}^{*}\to V_{m}$ (43)
@@
$\qquad=\int_{0}^{\pi}\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}\left(\delta\Omega\sin^{2}\theta\right)\right)\Psi\overline{z}\,d\theta=\int_{0}^{\pi}\frac{-1}{\sin\theta}\frac{d}{d\theta}\left(\delta\Omega\sin^{2}\theta\right)\frac{d}{d\theta}\left(\Psi\overline{z}\right)\,d\theta$
$\qquad=\int_{0}^{\pi}(\delta\Omega)(\theta)\sin^{2}\theta\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}\left(\Psi\overline{z}\right)\right)\,d\theta=\left\langle\delta\Omega\,,\,\frac{\sin}{r^{2}}\frac{d}{d\theta}\left(\frac{1}{\sin}\frac{d}{d\theta}\left(\overline{\Psi}z\right)\right)\right\rangle$
@@
$R^{*}:L^{2}(I^{\prime},r^{2}\sin)\to L^{2}(I,r^{2}\sin)\hskip 18.49988ptR^{*}\Psi)(\theta):=\begin{cases}\Psi(\theta),&\theta\in I^{\prime},\\ 0,&\theta\in I\setminus I^{\prime}.\end{cases}$
The corresponding forward operator relates to that of full measurement via $F_{\mathrm{part}}[p]=RF_{\mathrm{full}}[p]$. Thus, $F_{\mathrm{part}}^{\prime}[p]^{*}=F_{\mathrm{full}}^{\prime}[p]^{*}R^{*}$ with $F_{\mathrm{full}}^{\prime}[p]^{*}$ as in Corollary 12.
@@ 5 Tangential cone condition
As discussed in Section 4, ill-posedness of our inverse problem necessitates regularization strategies for stable reconstruction. Gradient-based regularization methods, including Landweber, Newton-type schemes, and their variants, rely on three core elements: well-posedness of the forward operator $F$ (Section 3), the adjoint of the linearized operator (Section 4.2), and convergence guarantees. The third requirement, which forms the primary focus of this section, is ensured by structural assumptions on $F$ – particularly, condition on nonlinearity and uniform boundedness of its derivative.
We address this through the tangential cone condition (TCC), a celebrated criterion first introduced in [HNS95] that ensures local convergence of iterative regularization algorithms. Intuitively, if the forward map $F$ is excessively nonlinear, there is no general guarantee that gradient descent steps will remain within the vicinity of the true solution. The tangential cone condition offers a quantitative tool for assessing nonlinearity of $F$, particularly of compact operators, thereby enabling application of regularization methods to realistic problems. Notably, this condition also implies: characterization of the solution set via null space of the linearized operator, uniqueness of minimum-norm solutions, and local convergence of reconstruction sequences towards minimum-norm solutions [KalNeuSch08].
For a general model $F:X\ni p\mapsto y\in Y$, the TCC states that the following inequality holds:
@@
where $C_{tc}>0$ is the TCC constant, and $B^{X}_{R}(p^{\dagger})$ is the ball of radius $R$ around a ground truth $p^{\dagger}\in\mathcal{D}(F)$. In our setting, the TCC reads as
@@
Weak variants of the TCC are discussed in [Kindermann17]. Beyond the TCC, other structural conditions can also ensure convergence guarantees: range-invariance [Kaltenbacher23-rangeinv], convexity, and the Polyak–Łojasiewicz condition [Nguyen24]
Although the TCC is a powerful tool, verification in practice can be challenging, even for specific examples. For a general verification strategy with full data, we refer the reader to [TCC21]. Applications of TCC verification span a wide range: elliptic inverse problems [HoffmanWaldNguyen:2021], elastography [Nakamura:MRE21, HubmerScherzer:TCC18], electrical impedance tomography [Kindermann21], full waveform inversion [EllerRolandRieder24], and neural network-based inverse problems [ScherzerHofmannNashed, AarsetHollerNguyen23].
Our contribution is the TCC for inertial wave inversion under full measurement by introducing a lifted regularity strategy (Section 3.4) that handles realistic $L^{2}$-data, and extending the verification to the restricted measurement regime, which is novel in this context. As a consequence, we obtain local unique identifiability of the unknown parameters, a property well-known to be important in ill-posed inverse problems.